OpenStudy (smurfy14):

√(115/12) Show steps please!!

OpenStudy (saifoo.khan):

\[\sqrt{\frac{115}{12}}\]

OpenStudy (saifoo.khan):

that?

OpenStudy (smurfy14):

yes

OpenStudy (anonymous):

√(115/12) =sqrt(115)/2sqrt(3) =3.095